# Channels of Stellar-mass Black Hole Formation

### v1

@@ 4.4 Channel 4: Black Hole Formation in the Context of No Supernova Explosion: The “Silent” Channel
Figure 2 shows the behavior of the shock radii for these representives of “silent” supernovae. In particular, the right hand panel indicates the spiral SASI motion quite clearly as the PNS and shock radii shrink. Figure 10 indicates that their early neutrino emissions are intermediate, as their compactnesses would suggest. Not shown are their late-time neutrino emissions, which due to continued accretion, should eventually overtake those of the exploding models. Figure 11 depicts the corresponding gravitational wave signatures, in comparison to those of the other models. We note that for these exemplars of the “silent” channel there is no matter memory effect, but that the gravitational-wave signal strength is dwarfed by those for the other black hole formation channels. The spiral SASI for the 12.25- and 14-$M_{\odot}$ models continues for the duration of the Fornax calculations (shown), but should continue much longer (not shown) and will modulate both the gravitational-wave and neutrino signals. Hence, as Figures 10 and 11 demonstrate, both their late-time gravitational-wave and neutrino emissions are in principle diagnostics of this modality of black hole formation.
@@ 5 Discussion and Conclusions
What emerges from a comparison of these channels of black hole formation is an interesting and provocative possible systematic trend. Channel 2 might give birth to black holes with masses from the lower gap to $\sim$10 $M_{\odot}$ in a weak to modest supernova explosion with a small kick that might range from zero to $\sim$100 km s-1. Such an event may be difficult to distinguish from a “regular” supernova. For this channel the energy and black hole mass may be anti-correlated and this might be, but has not been proven to be, a major modality for stellar-mass black hole formation. We emphasize that all the residual black hole and ejecta mass estimates are contingent upon the degree of pre-dynamical mass loss, either in winds or due to binary mass transfer. Both these effects can be significant and will be important factors in the resulting black hole birth mass function.
In the Channel 1 context, we see high-energy, asymmetrical supernova explosions that leave lower-gap and above black hole masses, but with high kick speeds that suggest they may not stay bound to a companion. Hence, though such an event might leave behind a lower-gap mass black hole, it might not occupy it observationally. The Channel 4 context is “silent” in that there is no supernova explosion, though there might be a secondary display, and the mass of the final black hole may more closely reflect the mass of the progenitor at collapse. The kicks for this channel due to supernova processes alone are due to net asymmetries in the neutrino emission and are likely to be low. However, they too could range from zero to perhaps $\sim$100 km s-1.
@@
Then, the question is whether the time between the inauguration of the supernova and the formation of a black hole at the associated high heating rates is adequate for the integral of the heating rate over time to result in a vigorous enough explosion to unbind most of the outer star. For our solar-metallicity 40-$M_{\odot}$ and 19.56-$M_{\odot}$ models, it was. For our tenth-solar 100-$M_{\odot}$ model, it wasn’t. The results were, on the one hand, lower gap and above black holes and strong, asymmetrical explosions and on the other no CCSN explosion (though in that case a PPISN explosion) and a 37-$M_{\odot}$ black hole perhaps just below the upper mass gap.
@@
Hence, Channel 1 black holes would seem to result in an interval, however narrow, of high compactnesses and mantle binding energies. Above this compactness range, a core-collapse supernova is launched, but aborted prematurely, and a much more massive black hole forms (Channel 3). Some of these may be PPISNe. Since lower-metallicity progenitors have higher compactnesses, for them the gap-forming compactness range may be narrower and the non-exploding, more massive black-hole formation channel may be wider. Therefore, whether a black hole is formed in one or the other of these channels depends upon metallicity and progenitor mass, and this highlights the centrality of stellar and binary evolution theory, still very much in flux. We note that the behavior and outcome of our 100-$M_{\odot}$ simulation, in particular vis à vis a possible metallicity dependence and the dissipation of the shock wave into a sound wave was anticipated in the earlier papers of Moriya et al. (2019), Chan et al. (2018), and Rahman et al. (2022).

### v2

@@ 4.4 Channel 4: Black Hole Formation in the Context of No Supernova Explosion: The “Silent” Channel
Figure 2 shows the behavior of the shock radii for these representives of “silent” supernovae. In particular, the right hand panel indicates the spiral SASI motion quite clearly as the PNS and shock radii shrink. Figure 11 indicates that their early neutrino emissions are intermediate, as their compactnesses would suggest. Not shown are their late-time neutrino emissions, which due to continued accretion, should eventually overtake those of the exploding models. Figure 12 depicts the corresponding gravitational wave signatures, in comparison to those of the other models. We note that for these exemplars of the “silent” channel there is no matter memory effect, but that the gravitational-wave signal strength is dwarfed by those for the other black hole formation channels. The spiral SASI for the 12.25- and 14-$M_{\odot}$ models continues for the duration of the Fornax calculations (shown), but should continue much longer (not shown) and will modulate both the gravitational-wave and neutrino signals. Hence, as Figures 11 and 12 demonstrate, both their late-time gravitational-wave and neutrino emissions are in principle diagnostics of this modality of black hole formation.
@@ 5 Discussion and Conclusions
What emerges from a comparison of these channels of black hole formation is an interesting and provocative possible systematic trend. We speculate that Channel 2 might give birth to black holes with masses from $\sim$3.0 $M_{\odot}$ to $\sim$10 $M_{\odot}$ in a weak to modest supernova explosion with a small kick that might range from zero to $\sim$100 km s-1. Such an event may be difficult to distinguish from a “regular” supernova. For this channel the energy and black hole mass may be anti-correlated and this might be, but has not been proven to be, a major modality for stellar-mass black hole formation. We emphasize that all the residual black hole and ejecta mass estimates are contingent upon the degree of pre-dynamical mass loss, either in winds or due to binary mass transfer. Both these effects can be significant and will be important factors in the resulting black hole birth mass function.
In the Channel 1 context, we see high-energy, asymmetrical supernova explosions that leave black holes, but with high kick speeds that suggest they may not stay bound to a companion. The Channel 4 context is “silent” in that there is no supernova explosion, though there might be a secondary display, and the mass of the final black hole may more closely reflect the mass of the progenitor at collapse. The kicks for this channel due to supernova processes alone are due to net asymmetries in the neutrino emission and are likely to be low. However, they too could range from zero to perhaps $\sim$100 km s-1.
@@
Then, the question is whether the time between the inauguration of the supernova and the formation of a black hole at the associated high heating rates is adequate for the integral of the heating rate over time to result in a vigorous enough explosion to unbind most of the outer star. For our solar-metallicity 40-$M_{\odot}$ and 19.56-$M_{\odot}$ models, it was. For our tenth-solar 100-$M_{\odot}$ model, it wasn’t. The results were, on the one hand, lower mass black holes (Shao, 2022) born in strong, asymmetrical explosions and on the other no CCSN explosion (though in that case a PPISN explosion) and a 37-$M_{\odot}$ black hole.
@@
Hence, Channel 1 black holes would seem to result in an interval, however narrow, of high compactnesses and mantle binding energies. Above this compactness range, a core-collapse supernova is launched, but aborted prematurely, and a much more massive black hole forms (Channel 3). Some of these may be PPISNe. Lower-metallicity progenitors have higher compactnesses. Hence, we speculate that whether a black hole is formed in one or the other of these channels depends upon metallicity and progenitor mass, and this highlights the centrality of stellar and binary evolution theory, still very much in flux. We note that the behavior and outcome of our 100-$M_{\odot}$ simulation, in particular vis à vis a possible metallicity dependence and the dissipation of the shock wave into a sound wave was anticipated in the earlier papers of Moriya et al. (2019), Chan et al. (2018), and Rahman et al. (2022).